Find the pH of the two equivalence points and the volume (mL) of 0.367M HNO3 needed to reach them in the titration of 0.188L of 0.250M ethylenediamine.

Find the pH of the two equivalence points and the
Answer #1

V1S1 = V2S2

V1 =? for HNO3

S1 = 0.367 M

V2 = 0.188 L for ethylenedimine

S2 = 0.250 M

Now, at equivalence point, the balanced equation is-

2HNO3 + (CH2)2(NH)2 = (CH2)2(NH3+)2][2NO3-]

2 equivalents of HNO3 required for complete neutralization.

So, volume HNO3 required for the equivalence point = 0.188*0.250/(0.367/2) L = 0.256 L

Total volume = 0.188+0.256 =0.444 L

So, final concentration of ethylenediamine = 0.188*0.250/0.444 M =0.106

pKb1 and pKb2 are required for finding the pH